Consider a relation R on a set A = { 2, 4, 7 }.



Given the relation R = { (2, 2), (2, 4), (2, 7), (4, 7}. Find:



1. Complement of a Relation



_____________________________________________________



_____________________________________________________





2. Inverse of a Relation



__________________________________________________________________________________________________________





3. Composite Product R o S and S o R ; S = { (1, 2), (2, 4), (2, 7) }

1
Expert's answer
2022-03-01T13:08:46-0500

Given "R = \{(2, 2), (2, 4), (2, 7), (4, 7)\}" is a relation on the set "A = \{2,4,7\}".


A relation on a set A is a subset of the cartesian product "A \times A".

Here, "A \times A = \{(2,2),(2,4), (2,7), (4,2), (4,4), (4,7), (7,2), (7,4),(7,7)\}"

1. Complement of R = The set of all elements in "A \times A" but not in "R" "= \{(4,2), (4,4), (7,2), (7,4), (7,7)\}"


2. Inverse of R = "R^{-1} = \{(b,a)\mid (a,b)\in R\}=\{(2,2), (4,2),(7,2), (7,4)\}"


3. Given, "S = \{ (1, 2), (2, 4), (2, 7) \}"

"R \circ S = \{(1,2), (1,4), (1,7), (2,7)\}\\
S \circ R = \{(2,4), (2,7)\}"